OpenStudy (kearsten):

fan and medal plz include steps number of real-number solutions 0=4x^2+9 i tried mathway but couldnt find the topic

OpenStudy (kearsten):

@nincompoop

OpenStudy (nincompoop):

\(4x^2+9 = 0 \rightarrow 4x^2 + 9 \color{red}{-9} = 0 \color{red}{-9} \) \(4x^2 = -9 \) then \(\large \frac{4x^2}{\color{red}{4}} = \frac{-9}{\color{red}{4}} \rightarrow x^2 = \frac{-9}{4}\) solve for x by obtaining the square root of both sides it will give you the solutions

OpenStudy (anonymous):

That seems right

OpenStudy (kearsten):

i still dont get its confusing

OpenStudy (anonymous):

It seems that way but it's really not

OpenStudy (kearsten):

can you help me?

OpenStudy (kearsten):

@isaac4321

OpenStudy (anonymous):

He pretty much got it under control I'm terrible at explaining stuff